Use the information in the following adjusted trial balance for the Wilson Trucking Company. Account Title Debit Credit Cash $ 8,100 Accounts receivable 16,500 Office supplies 2,000 Trucks 183,000 Accumulated depreciation—Trucks $ 37,698 Land 75,000 Accounts payable 12,100 Interest payable 3,000 Long-term notes payable 52,000 K. Wilson, Capital 175,147 K. Wilson, Withdrawals 19,000 Trucking fees earned 123,000 Depreciation expense—Trucks 24,315 Salaries expense 57,687 Office supplies expense 6,765 Repairs expense—Trucks 10,578 Totals $ 402,945 $ 402,945 1. Calculate the current ratio. Assume that the industry average for the current ratio is 1.5.
Answer #1

Formula for Current Ratio

Current Ratio = Current Assets / Current Liabilities

Current Assets = Cash + Accounts Receivables + Office Supplies

Current Assets = 8,100 + 16,500 + 2,000

Current Assets = $ 26,600

Current Liabilities = Accounts Payables + Interest Payable

Current Liabilities = 12,100 + 3,000

Current Liabilities = $ 15,100

Current Ratio = 26,600 / 15,100

Current Ratio = 1.7615 Times

Current Ratio = 1.7615:1